Women’s heights have a mean of 63.6 inches and a standard deviation of 2.5 inches. What is the



probability that a woman will have a height of higher than 70 inches?

1
Expert's answer
2022-06-14T00:22:29-0400
"P(X>70)=1-P(Z\le\dfrac{70-63.6}{2.5})"

"=1-P(Z\le2.56)\approx0.0052"